OpenStudy (anonymous):

Find cos A if b=9,c=15.

OpenStudy (lgbasallote):

this sounds like cosine law....

OpenStudy (anonymous):

Is that the whole question?

OpenStudy (anonymous):

or is it a^2 +b^2=c^2 ??

OpenStudy (anonymous):

..........from the fig. we are getting a = 12 this means given triangle satisfy pythagoras theorem.....so angle opp. i.e. angle C= 90 degree. so using cosA =9/15